Beale Manufacturing Company has a beta of 1.7, and Foley Industries has a beta of 0.35. The required return on an index fund that holds the entire stock market is 9.5%. The risk-free rate of interest is 5.5%. By how much does Beale's required return exceed Foley's required return? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

Required return for:

Beale=5.5+(9.5-5.5)*1.7=12.3%

Foley=5.5+(9.5-5.5)*0.35=6.9%

Hence excess return=12.3-6.9

=5.4%